# How to Find the Exact Pareto Front for Multi-Objective MDPs?

### v1

@@ Abstract
Multi-objective Markov Decision Processes (MDPs) are receiving increasing attention, as real-world decision-making problems often involve conflicting objectives that cannot be addressed by a single-objective MDP. The Pareto front identifies the set of policies that cannot be dominated, providing a foundation for finding Pareto optimal solutions that can efficiently adapt to various preferences. However, finding the Pareto front is a highly challenging problem. Most existing methods either (i) rely on traversing the continuous preference space, which is impractical and results in approximations that are difficult to evaluate against the true Pareto front, or (ii) focus solely on deterministic Pareto optimal policies, from which there are no known techniques to characterize the full Pareto front. Moreover, finding the structure of the Pareto front itself remains unclear even in the context of dynamic programming, where the MDP is fully known in advance. In this work, we address the challenge of efficiently discovering the Pareto front. By investigating the geometric structure of the Pareto front in MO-MDP, we uncover a key property: the Pareto front is on the boundary of a convex polytope whose vertices all correspond to deterministic policies, and neighboring vertices of the Pareto front differ by only one state-action pair of the deterministic policy, almost surely. This insight transforms the global comparison across all policies into a localized search among deterministic policies that differ by only one state-action pair, drastically reducing the complexity of searching for the exact Pareto front. We develop an efficient algorithm that identifies the vertices of the Pareto front by solving a single-objective MDP only once and then traversing the edges of the Pareto front, making it more efficient than existing methods. Furthermore, the entire Pareto front can be found in $V$ iterations, where $V$ represents the number of vertices on the Pareto front. Our empirical studies demonstrate the effectiveness of our theoretical strategy in discovering the Pareto front efficiently.
@@ 1 Introduction
The Pareto front approximation methods can be broadly divided into two categories. The first straightforward approach derives the Pareto front by solving single-objective problems obtained from scalarizing the multi-objective MDP for each preference and combining the resulting Pareto optimal policies [7]. Although state-of-the-art methods in the field of MO-RL leverage the similarity between Q-functions or policy representations of Pareto optimal policies for nearby preferences to avoid solving the optimal policy from scratch for each preference, thus reducing complexity [12, 10, 5], these methods still require traversing the continuous preference space to approximate the Pareto front. Another Pareto front finding method iteratively finds the preference that improves the current Pareto front the most to avoid blind traversing the preference space [13, 9]. However, this method only identifies deterministic policies on the Pareto front, which does not capture the entire Pareto front. In 2D (2-objectives) cases, constructing the whole Pareto front from deterministic Pareto optimal policies simply connects adjacent deterministic points by a straight line. However, in general cases, it is not that simple. For example, in a 3D (3-objectives) case shown in Fig. 1, deriving the entire Pareto front (as in Fig. 1(a)) from the vertices corresponding to deterministic Pareto optimal policies (as in Fig. 1(b)) alone still requires significant effort. Therefore, even in scenarios where the MDP is fully known, deriving the entire Pareto front through dynamic programming remains complex. This raises a fundamental question:
@@ 2 Related work
In the literature on both multi-objective optimization and reinforcement learning, two categories of problems are commonly considered. The first assumes that the (sequential) decision maker has preset or deduced preferences over multiple objectives in advance, and the goal is to find a Pareto optimal solution that performs well under these given preferences [14, 15, 16, 17]. The methods addressing this problem in multi-objective MDPs are specifically referred to as single-policy methods [18, 19, 20, 6, 7]. The second category focuses on finding a set of (approximate) Pareto optimal solutions that are diverse enough to represent the full range of the Pareto front [21, 22, 23]. This allows the (sequential) decision maker to choose the most preferable solution afterward. Due to its flexibility in adapting to changing preferences, this approach has received significant attention. This section gives an overview of the methods to get the Pareto optimal front estimation in multi-objective MDP and MO-RL.
@@
As early as in [24], the dynamic programming-based method was proposed to find the optimal Pareto front for MO-MDP, though the size of the candidate non-stationary policy set grows exponentially with the time horizon. To efficiently find the vertices on the Pareto front, [9] introduced the Optimistic Linear Support (OLS) method, which solves a single-objective problem along the direction that improves the current set of candidate policies the most. However, OLS can only find vertices on the Pareto front and does not provide a method for constructing the entire Pareto front from these vertices, which remains non-trivial.
@@
Authors of [25] proposed Pareto Q-learning, which learns a set of Pareto optimal policies in the MO-RL setting. Building on this, subsequent work has investigated training a universal Pareto optimal policy set, where the policy network takes preference embedding as input and outputs the corresponding Pareto optimal policy [10, 26, 27, 5]. These Q-learning variants rely on the maximization of Q-function estimations over the action space, making them suitable only for discrete action spaces. Recent works have extended Q-learning with policy networks to handle continuous action spaces [28, 29]. While these Q-learning methods update the Q-function for each preference vector by leveraging the Q-functions of similar visited preference vectors thus avoiding the need to retrain from scratch, they still essentially require updating the Q-function for each visited preference vector. Consequently, they still need to explore the preference space to achieve a near-accurate Pareto front. This process becomes computationally inefficient and impractical as the reward dimension increases.
@@
There are also policy-based MORL algorithms designed to find Pareto optimal policies [30, 31, 6, 12]. A concept similar to our approach, which directly searches over the Pareto front, is the Pareto-following algorithm proposed by [12]. This algorithm employs a modified policy-gradient method that adjusts the gradient direction to follow the Pareto front. While it can reduce the complexity of finding the Pareto front by avoiding the need to converge to the optimal policy from scratch, it still requires multiple policy-gradient steps to identify even a nearby Pareto optimal policy. Furthermore, it cannot guarantee comprehensive coverage of the estimated policies, nor can it ensure discovering the true Pareto front.
@@
In addition to value-based and policy-based methods that extend from single-objective RL, there are also heuristic approaches for combining policies that have shown promising performance in specific settings. For instance, Rewarded Soup [2, 11] proposes learning optimal policies for individual preference criteria and then linearly interpolating these policies to combine their capabilities. While [2] demonstrated that rewarded soup is optimal when individual rewards are quadratic with respect to policy parameters, this quadratic reward condition is highly restrictive and does not apply to general MDP settings.
@@ 3 Pareto front searching on MO-MDP
We first provide an overview of the proposed efficient Pareto front searching algorithm, followed by an introduction to the basic concepts behind MO-MDP. Finally, we present the details of the proposed algorithm.
@@ 3.1 Algorithm Overview
By applying the above steps to all policies in the queue $Q$ until the queue is empty ensures that the neighboring policies of all possible deterministic policies on the Pareto front have been visited. Since all vertices and their incident faces are traversed by the proposed algorithm, it guarantees full exploration of all faces of the Pareto front.
As the algorithm is guaranteed to consistently search along the Pareto front, which will be shown in Section 4, the total number of iterations is the same as the number of deterministic Pareto optimal policies.
@@
3.2 Preliminaries on MO-MDP
We consider a discounted MO-MDP $(\mathcal{S},\mathcal{A},{\mathbf{P}},{\bm{r}},\gamma)$, where $\mathcal{S}$ and $\mathcal{A}$ are the sets of states and actions, with cardinalities $S$ and $A$, respectively. The transition kernel ${\mathbf{P}}$ defines the probability ${\mathbf{P}}(s^{\prime}|s,a)$ of transitioning from $s$ to $s^{\prime}$ given action $a$. The reward tensor ${\bm{r}}\in\mathbb{R}^{S\times A\times D}$ specifies the $D$-dimensional reward ${\bm{r}}(s,a)$ obtained by taking action $a$ in state $s$. $\gamma<1$ is the discount factor.
Define $\Pi$ as the set of stationary policies, where for any $\pi\in\Pi$, $\pi(a|s)$ represents the probability of selecting $a$ at state $s$. We consider an ergodic MDP, meaning that any state is reachable from any other state by following a valid policy. The assumption of MDP ergodicity is commonly used in RL literature [5, 32].
The value function with policy $\pi$ is denoted as ${\mathbf{V}}^{\pi}(s)\in\mathbb{R}^{d}$, which is written as
Given an initial state distribution $\mu>0$, which assumes that the probability of starting from any state $s$ is non-zero, we define the long-term return of policy $\pi$ as $J^{\pi}(\mu)=\mathbb{E}_{s\sim\mu(\cdot)}\left[{\mathbf{V}}^{\pi}(s)\right]$. Let ${\mathbb{J}}(\mu)$ denote the set of expected long-term returns for all stationary policies under the initial state distribution $\mu$, i.e., ${\mathbb{J}}(\mu)=\{J^{\pi}(\mu)|\pi\in\Pi\}$. For simplicity of notation, we will omit $\mu$ in the following sections.
Lemma 1.
${\mathbb{J}}$ is a closed convex polytope, and its vertices are achieved by deterministic policies.
@@ 3.3 Pareto front searching
Neighbour Search As detailed in 11 and 12 of Algorithm 1, for any deterministic Pareto optimal policy $\pi$ popped from $Q$, we first evaluate all policies in $\Pi_{1}(\pi)\setminus{\mathbb{Z}}$, which differ from $\pi$ by only one state-action pair and have not yet been visited (with the visited policies stored in ${\mathbb{Z}}$). Once the long-term returns of $\Pi_{1}(\pi)$ are known, we select the set of non-dominated policies, $\Pi_{1,\text{ND}}(\pi)$, the set of policies in $\Pi_{1}(\pi)$ that cannot be dominated by any other policies in $\Pi_{1}(\pi)$. Specifically, the PPrune algorithm from [13] efficiently identifies the set of non-dominated elements from a given set.
@@
Figure 3: Convex polytope (blue planes) and its Pareto front (red edge and plane)
@@
2:Output: Pareto optimal faces $\mathcal{F}$ and Pareto optimal vertices $\mathcal{V}$
3:Initialize $\mathcal{F}\leftarrow\varnothing$, $\mathcal{V}\leftarrow\varnothing$, $Q_{\text{face}}\leftarrow\varnothing$, and $w(G)\leftarrow\varnothing$ for any $G$.
@@
5:Initialization: For all $G\in{\mathbb{F}}(J^{\pi},\mathrm{Conv}(J^{\Pi_{1,\text{ND}}}\cup\{J^{\pi}\}))$, $Q_{\text{face}}.\text{push}(G)$ and $w(G)\leftarrow\text{normal vector of }G$.
6:while $Q_{\text{face}}$ is not empty do
7:     $H\leftarrow Q_{\text{face}}.\text{pop}()$.
8:     if there exist vertices of $H$ not in $\mathcal{F}$ then // $H\notin\mathcal{F}$
9:         if $H$ has a feasible solution to Eq. 1 then
[…]
11:         else if $\text{dim}(H)>1$ then
12:              Let ${\mathcal{S}}$ be the set of $\text{dim}(H)-1$-dimensional faces of $H$ that intersect $J^{\pi}$.
13:              for each $\tilde{H}\in{\mathcal{S}}$ do
14:                  $w(\tilde{H}).\text{add}(w(H))$.
15:              end for
@@ 4.1 Distance-one Property on Boundary of ${\mathbb{J}}$
We present a theorem stating that the endpoints of any edge on the boundary of ${\mathbb{J}}$ correspond to deterministic policies that differ by only one state-action pair, almost surely. Specifically, for a finite MDP $(\mathcal{S},\mathcal{A},{\mathbf{P}},{\bm{r}},\gamma)$, where ${\mathbf{P}}$ is the transition probability kernel and ${\bm{r}}\geq 0$ is the reward function, the set of ${\mathbf{P}}$ and ${\bm{r}}$ for which deterministic policies on the edge of the Pareto front differ by more than one state-action pair has Lebesgue measure zero. The proof of this theorem is provided in Appendix D,
@@
For any edge on the boundary of ${\mathbb{J}}$ connecting two deterministic policies, these policies differ by only one state-action pair, almost surely.
@@ 4.2 Sufficiency of Traversing Over Edges
Lemma 2 only shows that a Pareto optimal vertex is connected to an edge (1-dimensional) on the Pareto front. However, it does not guarantee that one of the neighboring facets of a Pareto optimal vertex must also lie on the Pareto front. A facet of ${\mathbb{J}}$ lies on the Pareto front only if its normal vector is strictly positive (which will be shown later in Lemma 3), meaning that all points on the facet are not dominated by any other points in ${\mathbb{J}}$. However, the normal vectors of all facets intersecting at a Pareto optimal vertex are not necessarily strictly positive. This observation is consistent with Fig. 3, where point A is a Pareto optimal vertex, but it only lies on a Pareto optimal edge AB, and all its incident faces are not Pareto optimal. This also justifies that constructing the Pareto front in a 2D case (with two reward objectives) is relatively simple, as we only need to connect the neighbor vertices. However, in higher dimensions, the Pareto front becomes more complex, and it is impossible to build the Pareto front purely from vertices.

### v2

@@ Abstract
Multi-Objective Markov Decision Processes (MO-MDPs) are receiving increasing attention, as real-world decision-making problems often involve conflicting objectives that cannot be addressed by a single-objective MDP. The Pareto front identifies the set of policies that cannot be dominated, providing a foundation for finding Pareto optimal solutions that can efficiently adapt to various preferences. However, finding the Pareto front is a highly challenging problem. Most existing methods either (i) rely on traversing the continuous preference space, which is impractical and results in approximations that are difficult to evaluate against the true Pareto front, or (ii) focus solely on deterministic Pareto optimal policies, from which there are no known techniques to characterize the full Pareto front. Moreover, finding the structure of the Pareto front itself remains unclear even in the context of dynamic programming, where the MDP is fully known in advance. In this work, we address the challenge of efficiently discovering the Pareto front, involving both deterministic and stochastic Pareto optimal policies. By investigating the geometric structure of the Pareto front in MO-MDPs, we uncover a key property: the Pareto front is on the boundary of a convex polytope whose vertices all correspond to deterministic policies, and neighboring vertices of the Pareto front differ by only one state-action pair of the deterministic policy, almost surely. This insight transforms the global comparison across all policies into a localized search among deterministic policies that differ by only one state-action pair, drastically reducing the complexity of searching for the exact Pareto front. We develop an efficient algorithm that identifies the vertices of the Pareto front by solving a single-objective MDP only once and then traversing the edges of the Pareto front, making it more efficient than existing methods. Our empirical studies demonstrate the effectiveness of our theoretical strategy in discovering the Pareto front efficiently.
@@ 1 Introduction
The Pareto front approximation methods can be broadly divided into two categories. The first straightforward approach derives the Pareto front by solving single-objective problems obtained from scalarizing the multi-objective MDP for each preference and combining the resulting Pareto optimal policies [7]. Although state-of-the-art methods in the field of MO-RL leverage the similarity between Q-functions or policy representations of Pareto optimal policies for nearby preferences to avoid solving the optimal policy from scratch for each preference, thus reducing complexity [12, 10, 5], these methods still require sampling the continuous preference space to approximate the Pareto front. In cases where the Pareto front contains low-dimensional faces, sampling preference vectors on those faces makes it nearly impossible for approximation methods to identify such edge cases and accurately reconstruct the exact Pareto front. Another Pareto front finding method iteratively finds the preference that improves the current Pareto front the most to avoid blind traversing the preference space [13, 9]. However, this method only identifies deterministic policies on the Pareto front, which does not capture the entire Pareto front. In 2D (2-objectives) cases, constructing the whole Pareto front from deterministic Pareto optimal policies simply connects adjacent deterministic points by a straight line. However, in general cases, it is not that simple. For example, in a 3D (3-objectives) case shown in Fig. 1, deriving the entire Pareto front (as in Fig. 1(a)) from the vertices corresponding to deterministic Pareto optimal policies (as in Fig. 1(b)) alone still requires significant effort. Therefore, even in scenarios where the MDP is fully known, deriving the entire Pareto front through dynamic programming remains complex. This raises a fundamental question:
@@ 2 Related work
In the literature on both multi-objective optimization and reinforcement learning, two categories of problems are commonly considered. The first assumes that the (sequential) decision maker has preset or deduced preferences over multiple objectives in advance, and the goal is to find a Pareto optimal solution that performs well under these given preferences [14, 15, 16, 17]. The methods addressing this problem in multi-objective MDPs are specifically referred to as single-policy methods [18, 19, 20, 6, 7]. The second category focuses on finding a set of (approximate) Pareto optimal solutions that are diverse enough to represent the full range of the Pareto front [21, 22, 23]. This allows the (sequential) decision maker to choose the most preferable solution afterward. Due to its flexibility in adapting to changing preferences, this approach has received significant attention. However, finding the Pareto optimal front in the multi-objective optimization field does not take advantage of the specific structure of MO-MDPs, where multiple objectives share the same state and action spaces, as well as identical transition dynamics. Exploiting this structure in MO-MDPs could potentially lead to significant reductions in computational complexity. This section overviews the methods to get the Pareto optimal front estimation in MO-MDPs and MO-RL.
@@
As early as [24], the dynamic programming-based method was proposed to find the optimal Pareto front for MO-MDPs, though the size of the candidate non-stationary policy set grows exponentially with the time horizon. To efficiently find the vertices on the Pareto front, [9] introduced the Optimistic Linear Support (OLS) method, which solves a single-objective problem along the direction that improves the current set of candidate policies the most. However, OLS can only find vertices on the Pareto front and does not provide a method for constructing the entire Pareto front from these vertices, which remains non-trivial.
@@
[25] proposed Pareto Q-learning, which learns a set of Pareto optimal policies in the MO-RL setting. Building on this, subsequent work has investigated training a universal Pareto optimal policy set, where the policy network takes preference embedding as input and outputs the corresponding Pareto optimal policy [10, 26, 27, 5]. These Q-learning variants rely on the maximization of Q-function estimations over the action space, making them suitable only for discrete action spaces. Recent works have extended Q-learning with policy networks to handle continuous action spaces [28, 29]. While these Q-learning methods update the Q-function for each preference vector by leveraging the Q-functions of similar visited preference vectors thus avoiding the need to retrain from scratch, they essentially require updating the Q-function for each visited preference vector. Consequently, they still need to explore the preference space to achieve a near-accurate Pareto front. This process becomes computationally inefficient and impractical as the reward dimension increases.
@@
There are also policy-based MORL algorithms designed to find Pareto optimal policies [30, 31, 6, 12]. A concept similar to our approach, which directly searches over the Pareto front, is the Pareto-following algorithm proposed by [12]. This algorithm employs a modified policy-gradient method that adjusts the gradient direction to follow the Pareto front. While it can reduce the complexity of finding the Pareto front by avoiding the need to converge to the optimal policy from scratch, it still requires multiple policy-gradient stpng to identify even a nearby Pareto optimal policy. Furthermore, it cannot guarantee comprehensive coverage of the estimated policies, nor can it ensure discovering the true Pareto front.
@@
In addition to value-based and policy-based methods that extend from single-objective RL, there are also heuristic approaches for combining policies that have shown promising performance in specific settings. For instance, rewarded soup [2, 11] proposes learning optimal policies for individual preference criteria and then linearly interpolating these policies to combine their capabilities. While [2] demonstrated that rewarded soup is optimal when individual rewards are quadratic with respect to policy parameters, this quadratic reward condition is highly restrictive and does not apply to general MDPs.
@@ 3 Pareto front searching on MO-MDP
We begin with an introduction to the basic concepts underlying MO-MDPs. Next, we provide an overview of the proposed efficient Pareto front searching algorithm to offer intuition into its workings. Finally, we present a detailed explanation of the algorithm. A proof sketch of the proposed Pareto front searching algorithm will be provided in Section 4.
3.1 Preliminaries on MO-MDP
We consider a discounted finite MO-MDP $(\mathcal{S},\mathcal{A},{\mathbf{P}},{\bm{r}},\gamma)$, where $\mathcal{S}$ and $\mathcal{A}$ are the sets of states and actions, with cardinalities $S$ and $A$, respectively. The transition kernel ${\mathbf{P}}$ defines the probability ${\mathbf{P}}(s^{\prime}|s,a)$ of transitioning from $s$ to $s^{\prime}$ given action $a$. The reward tensor ${\bm{r}}\in\mathbb{R}^{S\times A\times D}$ specifies the $D$-dimensional reward ${\bm{r}}(s,a)$ obtained by taking action $a$ in state $s$. $\gamma<1$ is the discount factor.
Assumption 1.
(Sufficient Coverage of the Initial State Distribution) We assume that the initial state distribution $\mu>0$, meaning that the probability of starting from any state is larger than zero.
The sufficient coverage of the initial state distribution assumption is commonly used in RL [32]. Define $\Pi$ as the set of stationary policies, where for any $\pi\in\Pi$, $\pi(a|s)$ represents the probability of selecting $a$ at state $s$. The value function with policy $\pi$ is denoted as ${\mathbf{V}}^{\pi}(s)\in\mathbb{R}^{d}$, which is written as
We define the long-term return of policy $\pi$ as $J^{\pi}(\mu)=\mathbb{E}_{s\sim\mu(\cdot)}\left[{\mathbf{V}}^{\pi}(s)\right]$. Let ${\mathbb{J}}(\mu)$ denote the set of expected long-term returns for all stationary policies under the initial state distribution $\mu$, i.e., ${\mathbb{J}}(\mu)=\{J^{\pi}(\mu)|\pi\in\Pi\}$. For simplicity of notation, we will omit $\mu$ in the following sections.
Lemma 1.
In discounted finite MO-MDP, under 1, ${\mathbb{J}}$ is a closed convex polytope, and its vertices are achieved by deterministic policies.
@@ 3.2 Algorithm Overview
Applying the above steps to all policies in the queue $Q$ until the queue is empty ensures that the neighboring policies of all possible deterministic policies on the Pareto front have been visited. Since all vertices and their incident faces are traversed by the proposed algorithm, it guarantees full exploration of all faces of the Pareto front.
The intuition behind the proposed algorithm is that any deterministic Pareto optimal policy can identify all neighboring policies on the Pareto front by considering deterministic policies that differ by only one state-action pair. This ensures that all Pareto optimal policies are discovered during the search process. Additionally, the Pareto optimality justification excludes any policies that are not Pareto optimal, guaranteeing that the search trajectory remains on the Pareto front. The detailed theoretical results will be shown in Section 4. As the algorithm is guaranteed to consistently search along the Pareto front, the total number of iterations is the same as the number of deterministic Pareto optimal policies.
@@ 3.3 Pareto front searching
Neighbour Search As detailed in 11 and 12 of Algorithm 1, for any deterministic Pareto optimal policy $\pi$ popped from $Q$, we first evaluate all policies in $\Pi_{1}(\pi)\setminus{\mathbb{Z}}$, which differ from $\pi$ by only one state-action pair and have not yet been visited (with the visited policies stored in ${\mathbb{Z}}$). Once the long-term returns of $\Pi_{1}(\pi)$ are known, we select the set of non-dominated policies, $\Pi_{1,\text{ND}}(\pi)$, the set of policies in $\Pi_{1}(\pi)$ that cannot be dominated by any other policies in $\Pi_{1}(\pi)$. Specifically, the PPrune algorithm from [13] efficiently identifies the set of non-dominated elements from a given set, whose details are shown in Algorithm 3 in the appendix.
@@
Figure 3: Convex polytope and its Pareto front (red edge and plane)
@@
2:Initialize the set of Pareto optimal faces $\mathcal{F}\leftarrow\varnothing$, the set of Pareto optimal vertices $\mathcal{V}\leftarrow\varnothing$, the queue of faces for checking Pareto optimality $Q_{\text{face}}\leftarrow\varnothing$.
3:For all $G\in{\mathbb{F}}(J^{\pi},\mathrm{Conv}(J^{\Pi_{1,\text{ND}}}\cup\{J^{\pi}\}))$, $Q_{\text{face}}.\text{push}(G)$ and $w(G)\leftarrow\text{normal vector of }G$.
@@
5:while $Q_{\text{face}}$ is not empty do
6:     $H\leftarrow Q_{\text{face}}.\text{pop}()$.
7:     if there exist vertices of $H$ not in $\mathcal{F}$ then // $H\notin\mathcal{F}$
8:         if $w(H)$ has a feasible solution to Eq. 1 then
9:              .$\mathcal{F}.\text{add}(H)$ and $\mathcal{V}.\text{add}(\text{vertices of }H)$.
10:         else if $\text{dim}(H)>1$ then
11:              Let ${\mathcal{S}}$ be the set of $(\text{dim}(H)-1)$-dimensional faces of $H$ that intersect $J^{\pi}$.
12:              for each $\tilde{H}\in{\mathcal{S}}$ do
13:                  $Q_{\text{face}}.\text{push}(\tilde{H})$ and $w(\tilde{H}).\text{add}(w(H))$.
14:              end for
15:         end if
@@ 4.1 Distance-one Property on Boundary of ${\mathbb{J}}$
We present a theorem stating that the endpoints of any edge on the boundary of ${\mathbb{J}}$ correspond to deterministic policies that differ by only one state-action pair, almost surely. Specifically, for a finite MDP $(\mathcal{S},\mathcal{A},{\mathbf{P}},{\bm{r}},\gamma)$, where ${\mathbf{P}}$ is the transition probability kernel and ${\bm{r}}\geq 0$ is the reward function, the set of ${\mathbf{P}}$ and ${\bm{r}}$ for which deterministic policies on the edge of the Pareto front differ by more than one state-action pair has Lebesgue measure zero. The proof of this theorem is provided in Appendix D.
@@
In a discounted finite MO-MDP $(\mathcal{S},\mathcal{A},{\mathbf{P}},{\bm{r}},\gamma)$ and under 1, any edge on the boundary of ${\mathbb{J}}$ connecting two deterministic policies corresponds to policies that differ by only one state-action pair, almost surely.
@@ 4.2 Sufficiency of Traversing Over Edges
Lemma 2 only shows that a Pareto optimal vertex is connected to an edge (1-dimensional) on the Pareto front. However, it does not guarantee that one of the neighboring facets of a Pareto optimal vertex must also lie on the Pareto front. A facet of ${\mathbb{J}}$ lies on the Pareto front if and only if its normal vector is strictly positive (which will be shown later in Lemma 3), meaning that all points on the facet are not dominated by any other points in ${\mathbb{J}}$. However, the normal vectors of all facets intersecting at a Pareto optimal vertex are not necessarily strictly positive. This observation is consistent with Fig. 3, where point A is a Pareto optimal vertex, but it only lies on a Pareto optimal edge AB, and all its incident faces are not Pareto optimal. This also justifies that constructing the Pareto front in a 2D case (with two reward objectives) is relatively simple, as we only need to connect the neighbor vertices. However, in higher dimensions, the Pareto front becomes more complex, and it is impossible to build the Pareto front purely from vertices.